# In-Context Learning as Nonparametric Conditional Probability Estimation: Risk Bounds and Optimality

### v1

@@ 5 Lower Bounds
To establish the optimality of the convergence rate in Theorem 15, we derive a matching minimax lower bound for the ICL risk. Our argument leverages the classical minimax theory for nonparametric estimation under the squared Hellinger distance translate these results to the KL divergence setting.
The squared Hellinger distance, defined as $H^{2}(P,Q)=\frac{1}{2}\int(\sqrt{dP}-\sqrt{dQ})^{2}$, is a fundamental tool for deriving statistical lower bounds due to its favorable metric properties (van der Vaart and Wellner, 1996; Geer, 2000). The rate $T^{-\frac{2\beta}{2\beta+d}}$ is known to be minimax-optimal for estimation under the squared Hellinger loss (Tsybakov, 2008, Chapter 2). The key to translating this result is the classical inequality $2H^{2}(P,Q)\leq\text{KL}(P,Q)$, which holds under bounded likelihood ratio conditions, see (7.5) in Birgé and Massart (1998) and Lemma 4 in Haussler and Opper (1997). Conversely, under some integrability assumptions, the KL divergence can be bounded by the Hellinger distance up to logarithmic factors (Wong and Shen, 1995, Theorem 5). This two-sided relationship enables the direct translation of minimax lower bounds from the Hellinger distance to the KL divergence. The following lemma formalizes this connection for the truncated KL divergence.
@@
Suppose $B\geq 2$. Then the squared Hellinger distance and the truncated KL divergence satisfy
$H^{2}(P,Q)\leq\tfrac{1}{2}\mathrm{KL}_{2}(P,Q)\leq\tfrac{1}{2}\mathrm{KL}_{B}(P,Q)\leq 5e^{B/2}H^{2}(P,Q).$
@@
Consider the same setting as in Theorem 15. Then, the minimax lower bound of the ICL risk is given by
$\inf_{\hat{\mathbf{p}}}\sup_{\mathbf{p}_{0}\in\mathcal{G}_{1}(\beta,Q)}R_{B}(\mathbf{p}_{0},\hat{\mathbf{p}})\gtrsim T^{-\frac{2\beta}{2\beta+d}}.$
This result aligns with the upper bound in Theorem 15 up to constant factors, thereby establishing the minimax-optimality of the rate $T^{-\frac{2\beta}{2\beta+d}}$ for ICL under the truncated KL loss when $\alpha=1$. The case of general $\alpha\in(0,1)$ remains open and will be studied in future work.
@@ 6 In-Context Learning Beyond Transformers
ICL is often regarded as a distinctive capability of transformer architectures, with explanations frequently grounded in constructions based on attention mechanisms (Akyürek et al., 2023; Von Oswald et al., 2023; Zhang et al., 2024; Reddy, 2024; Lu et al., 2024).
However, recent literature suggests that ICL is not exclusive to transformer architectures. Tong and Pehlevan (2025) empirically demonstrated that MLPs can also learn in-context on standard synthetic ICL taks. Similarly, Abedsoltan et al. (2024) showed that standard MLPs are capable of task-scaling in ICL settings. Moreover, Kratsios and Furuya (2025) theoretically established that MLPs possess universal approximation capabilities for ICL tasks.
In this section, we support these insights by analyzing the expected excess risk of MLP-based classifiers in the setting of multiclass ICL. Under appropriate assumptions, we demonstrate that MLPs can achieve minimax-optimal convergence rates for ICL. Our analysis of convergence rates provides theoretical support for recent empirical evidence on the effectiveness of MLPs. This finding suggests that, while attention mechanisms are powerful, they are not strictly necessary for enabling ICL.
Specifically, we consider a class of MLPs with ReLU activation functions, used as probabilistic classifiers. To accommodate input prompts in $\mathbb{R}^{d}$, we modify the network architecture introduced in (4) by increasing its input dimension from $K$ to $d$. Accordingly, we introduce a function class as follows:
@@
where $\boldsymbol{\Phi}$ denotes the softmax function.
@@
For every $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ and $0<\varepsilon<\min\{(K(4+C_{Q,\beta,d}))^{-1},(\beta+1)^{-\beta},(Q+1)^{-\beta/d}e^{-\beta}\}$ with $C_{Q,\beta,d}=(2Q+1)(1+d^{2}+\beta^{2})6^{d}+Q3^{\beta}$, there exist a neural network $\tilde{\mathbf{p}}\in\mathcal{F}_{\rm MLP}(L,W,S)$ with
@@
$\tilde{p}_{k}\geq\frac{\varepsilon}{4},\quad\|\tilde{p}_{k}-p_{k}^{0}\|_{\infty}\leq\frac{1}{2}K(C_{Q,\beta,d}+4)\varepsilon,$
for all $k=1,\dots,K$.
This result slightly modifies Lemma 4.3 in Bos and Schmidt-Hieber (2022), which builds upon a line of work showing that ReLU MLPs can efficiently approximate $\beta$-Hölder smooth functions, with $L\lesssim\log(1/\varepsilon)$, $W\lesssim\varepsilon^{-d/\beta}$, $s\lesssim\varepsilon^{-d/\beta}\log(1/\varepsilon)$ (Yarotsky, 2017; Petersen and Voigtlaender, 2018; Schmidt-Hieber, 2020; Jiao et al., 2023; Chakraborty and Bartlett, 2024b).
@@
Consider the function class $\mathcal{F}_{\mathrm{MLP}}(L,W,S)$ defined in (9). For any $T\in\mathbb{N}$, the empirical covering entropy of the function class $\log(\mathcal{F}^{\prime}_{\Phi})$ under the empirical sup-norm satisfies
$\mathcal{V}_{\infty}\left(\tfrac{1}{T},\log(\mathcal{F}_{\rm MLP}(L,W,S)),T\right)\leq C\cdot SL\log S\cdot\log T,$
for some universal constant $C>0$.
The proof of Lemma 19 is provided in Appendix H. Building on Lemmas 18 and 19, we establish an upper bound on the expected excess risk for MLP-based classifiers in the ICL setting, stated as follows.
@@
Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$. Let $\hat{\mathbf{p}}$ be an estimator of $\mathbf{p}_{0}$, taking values in the function class $\mathcal{F}_{\rm MLP}(L,W,S)$, where the parameters of the class scale with the precision parameter $\varepsilon>0$ as follows:
@@
[…]
and we choose the precision parameter $\varepsilon$ as
$\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}},$
for some constant $C>0$, then the ICL risk is bounded by
$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}K^{4+\alpha}\log^{4}\left(KT\right).$
@@
$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}\right)\lesssim B^{2}T^{-\frac{2\beta}{2\beta+d}}K^{5}\log^{4}\left(KT\right).$
@@
We first show that the risk bounds in Theorems 20 and 15 both achieve the optimal rate $T^{-\frac{2\beta}{2\beta+d}}$, up to logarithmic factors. Recent works have empirically shown that MLPs can also exhibit ICL behavior (Tong and Pehlevan, 2025; Kratsios and Furuya, 2025; Abedsoltan et al., 2024). Building on this, we prove that in the task-scaling regime, where the number of tasks $T$ increases while the context length remains fixed, the convergence rate of MLPs matches that of transformer-based models. This provides a theoretical justification for the comparable performance of MLP-based ICL under realistic pretraining conditions.
Proof of Theorem 20. The approximation error is controlled as shown in Lemma 18. Using these bounds, we construct an estimator $\hat{\mathbf{p}}\in\mathcal{F}_{MLP}(L,W,S)$ such that $\|\hat{\mathbf{p}}-\mathbf{p}_{0}\|_{\infty}\leq\varepsilon$. Applying Theorems 711, and Lemma 19, and expressing the network parameters as functions of $\varepsilon$, we obtain the claimed in-context risk bound. $\blacksquare$
@@ 7 Discussions
The upper bound established in Theorem 15 reflects the curse of dimensionality, as the input prompt dimension $d$ explicitly appears in the convergence rate exponent. This phenomenon is well-known in nonparametric estimation under smoothness constraints, where the optimal minimax rate for estimating a $\beta$-Hölder function scales as $T^{-2\beta/(2\beta+d)}$ (Donoho and Johnstone, 1998). Our result demonstrates that this dependence persists in transformer-based ICL under general smoothness assumptions.
Nevertheless, the practical impact of dimensionality may be mitigated, as the number of tasks $T$ is typically large in realistic ICL scenarios. Furthermore, transformer architectures are believed to possess an inherent ability to adapt to latent low-dimensional structure within the input, potentially reducing their effective dependence on the prompt dimension.
Recent work has explored multiple strategies to overcome the curse of dimensionality. One strategy is to explore alternative function classes with more favorable approximation properties, such as mixed-Besov spaces (Suzuki, 2019) and Korobov spaces (Montanelli and Du, 2019). Another approach is to impose structural assumptions on the target function, including additive ridge functions (Fang and Cheng, 2023), hierarchical compositions (Schmidt-Hieber, 2020; Han et al., 2020), generalized single-index models (Bauer and Kohler, 2019), and multivariate adaptive regression splines (MARS) (Kohler et al., 2022).
@@
Beyond structural assumptions on function classes, recent literature highlights the importance of leveraging the geometric structure of input data. The manifold hypothesis (Pope et al., 2021) suggests that although data appear in high-dimensional spaces, their intrinsic structure is governed by a low-dimensional manifold (Tenenbaum et al., 2000; Roweis and Saul, 2000). Under this assumption, deep neural networks have been shown to achieve significantly faster convergence rates by effectively adapting to the intrinsic dimensionality of data (Nakada and Imaizumi, 2020; Chen et al., 2022, 2019; Schmidt-Hieber, 2019; Cloninger and Klock, 2021). Recent studies further characterize the statistical properties of deep learning and generative models under low intrinsic dimensionality, providing refined generalization guarantees in these settings (Chakraborty and Bartlett, 2024a, b).
@@
Despite these advances, the theoretical understanding of transformers in low-dimensional geometric settings remains limited. A recent contribution by Havrilla and Liao (2024) analyzes transformer approximation for a regression problem on low-dimensional manifolds, whose setting is quite different from ours. Extending Lemma 8 to manifold-supported functions deserve further study, as this would allow risk bounds to scale with the intrinsic dimension rather than the ambient dimension.

### v2

@@ 5 Lower Bounds
To establish the optimality of the convergence rate in Theorem 15, we derive a matching minimax lower bound for the ICL risk. Our argument leverages the classical minimax theory for nonparametric estimation under the squared Hellinger distance and translates these results to the KL divergence setting. The squared Hellinger distance between two probability measures $P$ and $Q$ is defined as $H^{2}(P,Q)=\frac{1}{2}\int(\sqrt{dP}-\sqrt{dQ})^{2}$, which is a fundamental tool for deriving statistical lower bounds due to its favorable metric properties (van der Vaart and Wellner, 1996; van de Geer, 2000). The rate $T^{-2\beta/(2\beta+d)}$ is known to be minimax optimal for nonparametric estimation under the squared Hellinger loss (Tsybakov, 2008, Chapter 2). The key to translating this result is the classical inequality $2H^{2}(P,Q)\leq\text{KL}(P\|Q)$, see Lemma 4 in Haussler and Opper (1997) and (7.5) in Birgé and Massart (1998). Conversely, under certain integrability assumptions, the KL divergence can be bounded by the Hellinger distance up to logarithmic factors (Wong and Shen, 1995, Theorem 5). This two-sided relationship enables the direct translation of minimax lower bounds from the Hellinger distance to the KL divergence. The following lemma formalizes this connection for the truncated KL divergence.
@@
Let $P$ and $Q$ denote two probability measures defined on the same measurable space. Then for any $B\geq 2$,
$H^{2}(P,Q)\leq\tfrac{1}{2}\mathrm{KL}_{2}(P\|Q)\leq\tfrac{1}{2}\mathrm{KL}_{B}(P\|Q)\leq 5e^{B/2}H^{2}(P,Q).$
@@
Consider the same setting as in Theorem 15. Then the minimax lower bound of the ICL risk with $\alpha=1$ is given by
$\inf_{\hat{\mathbf{p}}}\sup_{\mathbf{p}_{0}\in\mathcal{G}_{1}(\beta,Q)}R_{B}(\mathbf{p}_{0},\hat{\mathbf{p}})\gtrsim T^{-\frac{2\beta}{2\beta+d}},$
where $\mathcal{G}_{\alpha}(\beta,Q)=\{\mathbf{p}_{0}:\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)\ {\rm and}\ \mathbf{p}_{0}\ is\ \alpha-SVB\}$.
This result aligns with the upper bound in Theorem 15 up to logarithmic factors, thereby establishing the minimax optimality of the convergence rate for the ICL risk under the truncated Kullback-Leibler loss when $\alpha=1$. For general $\alpha\in(0,1)$, the minimax optimal rate of the expected excess risk under the truncated KL divergence remains unknown. We conjecture that the optimal rate is given by $T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}$, up to logarithmic factors. However, a rigorous proof is beyond the scope of the current paper and warrants further study.
@@ 6 In-Context Learning Beyond Transformers
ICL is often regarded as a distinctive capability of transformer architectures, with explanations frequently grounded in constructions based on attention mechanisms (Akyürek et al., 2023; Von Oswald et al., 2023; Zhang et al., 2024; Reddy, 2024; Lu et al., 2024). However, recent literature suggests that ICL is not exclusive to transformer architectures. Tong and Pehlevan (2025) empirically demonstrated that MLPs can also learn in-context on standard synthetic ICL tasks. Similarly, Abedsoltan et al. (2024) showed that standard MLPs are capable of task-scaling in ICL settings. Moreover, Kratsios and Furuya (2025) theoretically established that MLPs possess universal approximation capabilities for ICL tasks.
In this section, we support these insights by analyzing the expected excess risk of MLP-based classifiers in the setting of multiclass ICL. Under appropriate assumptions, we demonstrate that MLPs can achieve the minimax optimal convergence rate for ICL. Our analysis provides theoretical support for recent empirical evidence on the effectiveness of MLPs. This finding suggests that, while attention mechanisms are powerful, they are not strictly necessary for enabling ICL. Specifically, we consider a class of MLPs with ReLU activation functions, used as probabilistic classifiers. To accommodate input prompts in $\mathbb{R}^{d}$, we modify the network architecture introduced in (4) by increasing its input dimension from $K$ to $d$. Accordingly, we introduce a function class as follows:
@@
where $\boldsymbol{\Phi}$ denotes the softmax function and $\operatorname{vec}(\mathbf{Z})$ is the vectorization of the matrix $\mathbf{Z}$.
@@
For any $\mathbf{p}_{0}\in\mathcal{G}(\beta,Q)$ and $\varepsilon\in(0,\min\{(K(4+C_{Q,\beta,d}))^{-1},(\beta+1)^{-\beta},(Q+1)^{-\beta/d}e^{-\beta}\})$ with $C_{Q,\beta,d}=(2Q+1)(1+d^{2}+\beta^{2})6^{d}+Q3^{\beta}$, there exist a neural network $\tilde{\mathbf{p}}\in\mathcal{F}_{\rm mlp}(L,W,S)$ with
@@
$\|\tilde{p}_{k}-p_{k}^{0}\|_{\infty}\leq\frac{1}{2}K(C_{Q,\beta,d}+4)\varepsilon,\quad\tilde{p}_{k}(\mathbf{Z})\geq\frac{\varepsilon}{4}$
for all $k\in[K]$ and $\mathbf{Z}\in[0,1]^{d}$.
This result is a slight modification of Lemma 4.3 in Bos and Schmidt-Hieber (2022), which builds upon a line of work (Yarotsky, 2017; Schmidt-Hieber, 2020; Jiao et al., 2023; Chakraborty and Bartlett, 2024) showing that ReLU MLPs with depth $L\lesssim\log(1/\varepsilon)$, width $W\lesssim\varepsilon^{-d/\beta}$, and size $S\lesssim\varepsilon^{-d/\beta}\log(1/\varepsilon)$ can efficiently approximate $\beta$-Hölder smooth functions. Building on Lemma 18, we then derive a non-asymptotic upper bound for the approximation error $\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})$.
Theorem 19
Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$ and $\mathbf{p}=(p_{1},\ldots,p_{K})^{\top}:[0,1]^{d}\to\mathcal{S}^{K}$ is an approximation of $\mathbf{p}_{0}$ such that $\|\mathbf{p}-\mathbf{p}_{0}\|_{\infty}\leq C_{0}\varepsilon/4$ and $\min_{k}\inf_{\mathbf{Z}\in[0,1]^{d}}p_{k}(\mathbf{Z})\geq\varepsilon/4$ with $C_{0}$ being a positive constant. Then we have
$R(\mathbf{p}_{0},\mathbf{p})\lesssim K(C_{0}+1)^{2+\alpha}\varepsilon^{1+\alpha}\left(1+\frac{\mathbf{1}(\alpha<1)}{1-\alpha}+\log\left(\frac{4}{\varepsilon}\right)\right).$
Furthermore, if we choose $C_{0}=2K(C_{Q,\beta,d}+4)$ and $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ for some constant $C>0$, then
$\inf_{\mathbf{p}\in\mathcal{F}}R(\mathbf{p}_{0},\mathbf{p})\lesssim K^{3+\alpha}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}\log(KT).$
The proof of this theorem follows the same arguments as in Theorem 10. A similar result can be found in Bos and Schmidt-Hieber (2022, Theorem 3.2).
Remark 20
Similar to the arguments in Remark 30, by adding additional ReLU layers to the network in $\mathcal{F}_{\rm id}(L,W,S)$, we can assume that every network $\mathbf{G}=(G_{1},\dots,G_{p})^{\top}\in\mathcal{F}_{\rm id}(L,W,S)$ satisfies $\log(\varepsilon)\leq G_{i}\leq\log(2)$ for all $1\leq i\leq p$. Since $\mathcal{F}_{\rm mlp}(L,W,S)=\mathbf{\Phi}(\mathcal{F}_{\rm id}(L,W,S))$, it follows that every network in $\mathcal{F}_{\rm mlp}(L,W,S)$ has a uniform, strictly positive lower bound of $\varepsilon/2K$ for all its coordinates. Given that the neural network $\tilde{p}_{k}\geq\varepsilon/4\geq\varepsilon/2K$, this implies that $\tilde{\mathbf{p}}\in\mathcal{F}_{\rm mlp}(L,W,S)$. Therefore, with a suitable chosen precision parameter $\varepsilon$, the function class $\mathcal{F}_{\rm mlp}(L,W,S)$ satisfies Assumption 2. This allows us to apply the oracle inequality to derive the expected excess risk for MLP-based classifiers in the ICL setting.
The next lemma establishes an explicit upper bound on the uniform empirical entropy of $\mathcal{F}_{\rm mlp}(L,W,S)$, a key quantity in deriving the generalization error bound in the oracle inequality. We write $\mathcal{F}_{\rm mlp}(L,W,S)=\mathbf{\Phi}(\mathcal{F}_{\mathrm{id}}(L,W,S))$.
@@
Suppose $\|\mathbf{G}\|_{\infty}\leq B_{2}<\infty$ for any $\mathbf{G}\in\mathcal{F}_{\mathrm{id}}(L,W,S)$. Then
$\mathcal{V}_{\infty}\left(\frac{1}{T},\log(\mathcal{F}_{\rm mlp}(L,W,S)),T\right)\leq CKSL\log(S)\log(B_{2}KT),$
where $\mathcal{V}_{\infty}\left(1/T,\log(\mathcal{F}_{\rm mlp}(L,W,S)),T\right)=\sup_{\mathcal{Z}_{T}}\log\mathcal{N}(1/T,\log(\mathcal{F}_{\rm mlp}(L,W,S)),\|\cdot\|_{\infty,\mathcal{Z}_{T}})$, and $C>0$ is an universal constant.
The proof of Lemma 21 is provided in Appendix H. Building on the oracle inequality (Theorem 7), Theorem 19, and Lemma 21, we derive an upper bound on the expected excess risk for MLP-based classifiers in the ICL setting, as stated below.
@@
Assume $\mathbf{p}_{0}\in\mathcal{G}_{\alpha}(\beta,Q)$ for some $\alpha\in[0,1]$. Let $\hat{\mathbf{p}}_{\rm mlp}$ be an estimator of $\mathbf{p}_{0}$, taking values in the function class $\mathcal{F}_{\rm mlp}(L,W,S)$, where the parameters of this class scale with the precision parameter $\varepsilon>0$ as follows:
@@
If we choose the precision parameter $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ for some positive constant $C$, and the expected empirical risk gap satisfies
$\Delta_{T}(\mathbf{p}_{0},\widehat{\mathbf{p}}_{\rm mlp})\lesssim K^{5+\alpha}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}},$
then the ICL risk is bounded by
$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}_{\rm mlp}\right)\lesssim B^{2}K^{5+\alpha}T^{-\frac{(1+\alpha)\beta}{(1+\alpha)\beta+d}}\log^{6}\left(KT\right).$
@@
$R_{B}\left(\mathbf{p}_{0},\hat{\mathbf{p}}_{\rm mlp}\right)\lesssim B^{2}K^{6}T^{-\frac{2\beta}{2\beta+d}}\log^{6}(KT).$
@@
We have shown that the ICL risk bounds in Theorems 15 and 22 both achieve the optimal rate $T^{-2\beta/(2\beta+d)}$, up to logarithmic factors. Recent works have empirically shown that MLPs can also exhibit ICL behavior (Tong and Pehlevan, 2025; Kratsios and Furuya, 2025; Abedsoltan et al., 2024). Building on this, we prove that in the task-scaling regime, where the number of tasks $T$ increases while the context length remains fixed, the convergence rate of MLPs matches that of transformer-based models. This provides a theoretical justification for the comparable performance of MLP-based ICL under realistic pretraining conditions.
Proof of Theorem 22. Recall that $\varepsilon=C^{-1}T^{-\frac{\beta}{(1+\alpha)\beta+d}}K^{-\frac{(3+\alpha)\beta}{(1+\alpha)\beta+d}}$ and $\|\mathbf{G}\|_{\infty}\leq|\log(\varepsilon)|$ for any $\mathbf{G}\in\mathcal{F}_{\mathrm{id}}(L,W,S)$, it follows that $\|\mathbf{G}\|_{\infty}\leq C\log(KT)$ for some positive constant $C$. Applying Theorems 719, Lemmas 18 and 21, and expressing the network parameters as functions of $\varepsilon$, we obtain the claimed in-context risk bound. $\blacksquare$
@@ 7 Discussions
The upper bounds established in Theorem 15 and  22 reflect the curse of dimensionality, as the input prompt dimension $d$ explicitly appears in the convergence rate exponent. This phenomenon is well-known in nonparametric estimation under smoothness constraints, where the optimal minimax rate for estimating a $\beta$-Hölder function scales as $T^{-2\beta/(2\beta+d)}$ (Donoho and Johnstone, 1998). Our result demonstrates that this dependence persists in transformer-based ICL under general smoothness assumptions. Nevertheless, the practical impact of dimensionality may be mitigated, as the number of tasks $T$ is typically extremely large in realistic ICL scenarios. Furthermore, transformer architectures are believed to possess an inherent ability to adapt to latent low-dimensional structure within the input, potentially reducing their effective dependence on the prompt dimension.
@@
Recent work has explored multiple strategies to overcome the curse of dimensionality. One strategy is to explore alternative function classes with more favorable approximation properties, such as mixed-Besov spaces (Suzuki, 2019) and Korobov spaces (Montanelli and Du, 2019). Another approach is to impose structural assumptions on the target function, including additive ridge functions (Fang and Cheng, 2023), hierarchical compositions (Schmidt-Hieber, 2020; Han et al., 2020), generalized single-index models (Bauer and Kohler, 2019), and multivariate adaptive regression splines (MARS) (Kohler et al., 2022). Beyond structural assumptions on function classes, recent literature highlights the importance of leveraging the geometric structure of input data. The manifold hypothesis (Pope et al., 2021) suggests that although data appear in high-dimensional spaces, their intrinsic structure is governed by a low-dimensional manifold (Tenenbaum et al., 2000; Roweis and Saul, 2000). Under this assumption, deep neural networks have been shown to achieve significantly faster convergence rates by effectively adapting to the intrinsic dimensionality of data (Schmidt-Hieber, 2019; Chen et al., 2019, 2022; Nakada and Imaizumi, 2020; Cloninger and Klock, 2021). Recent studies further characterize the statistical properties of deep learning and generative models under low intrinsic dimensionality, providing refined generalization guarantees in these settings (Chakraborty and Bartlett, 2024, 2025).
@@
Despite these advances, the theoretical understanding of transformers in low-dimensional geometric settings remains limited. A recent contribution by Havrilla and Liao (2024) analyzes transformer approximation for a regression problem on low-dimensional manifolds, whose setting is quite different from ours. Extending Lemma 8 to manifold-supported functions warrants further study, as this would allow risk bounds to scale with the intrinsic dimension rather than the ambient dimension.
Acknowledgments and Disclosure of Funding
We gratefully acknowledge the support of the National Natural Science Foundation of China (NSFC12471276, NSFC12131006, NSFC12201048, NSFC12301383, NSFC12471276) and the Scientific and Technological Innovation Project of China Academy of Chinese Medical Science (CI2023C063YLL).